# ArtA: Automating Design Space Exploration of Spin Qubit Architectures

### v2

@@ VII.2 Architecture Analysis
Table 5: Architectures returned by ArtA with BO [EI, MA] optimization method for a range of quantum circuits. Results are color-coded, with color schemes following the categories introduced in Section IV.
@@
In Table 5, we present the values of the architectural variables, introduced in Section IV, that were found by ArtA having the highest ESP with the BO [EI, MA] optimization method alongside their ESP values. These circuits are organized by class and size in number of qubits within the rows of the table. Subsequent columns present the ESP of the best architecture discovered, followed by color-coded architectural parameter categories as delineated in Section IV. The use of color accents in the table serves to enhance visual clarity, with each corresponding to the values of each variable.
@@
First, we observe a unanimous preference for the Z rotation implementation (z_rot_impl) in the zri column, with a pulse-based rotation being preferred. Note that previously, we have assumed in Section V.1 that a shuttle and a pulse-based Z rotation share the same single-qubit gate fidelity, but because the latter requires two shuttles, it can benefit the ESP more than the other type. Moreover, as a shuttle operation involves two quantum dots while a pulse-based operation only involves one, the shuttle-based Z rotation will have a higher chance of incurring crosstalk errors based on our crosstalk definition. Given these two reasons, BO has favored the pulsed-based version. This indicates that shuttle-based single-qubit rotations are not as sustainable as pulsed-based for large-scale circuits.
Moving on, since the X, Y, and Z gates are implemented the same way hardware-wise, xy_z automatically is assumed 1 based on the interdependencies detailed in Section IV.6. Consequently, xy_tqg and z_tqg are equivalent, taking both either 0 or 1 from the optimization method of ArtA (BO [EI, MA]), as we can observe in the table. Additionally, xy_z_tqg can also be conceptually equivalent to xy_tqg and z_tqg since the pulse-based Z rotation is selected. Therefore, whenever xy_z_tqg is 1 or xy_tqg and z_tqg is 1, it means the architecture(s) which allow(s) parallelization of all single- and two-qubit gates obtains the highest ESP. In that sense, one would expect xy_z_tqg, xy_tqg and z_tqg to be synchronized. However, since such exploration freedom was given in the design space, the reasons for the mismatches could be explained by the internal characteristics of the circuits themselves. For example, in QFT100, the parallelization of all gate types occurs more often than in QFT60, which has xy_z_tqg = 0, xy_tqg = 1, and z_tqg = 1.
Conversely, when applying the same reasoning to QV8, QV10, and QFT20, it seems unnecessary for the hardware to support parallel executions of single-qubit and two-qubit gates. This shows that the success of real quantum circuits can come from different “angles” with reduced hardware capabilities  [^3] Ultimately, each circuit profile has specific architectural traits that can be leveraged to achieve optimal performance without the need for excessively advanced hardware features (achieving a balance in architectural characteristics)..
@@
The same can be observed in our universal architecture recommendation in the last row, which performs best on average for all the circuits examined. Additionally, the operational fidelity, which is directly linked to the gate overhead of compilation, can also severely impact the final ESP for circuits with high gate counts, which in turn explains the need to add the least additional gates possible during compilation through the use of beSnake [50] and local single-qubit implementation. Having said that, Bernstein Vazirani, QV8, and QV10 circuits preferred the global implementation, as these circuits consist of a significant number of similar single-qubit gates being executed at the same time step.
@@
Overall, we can observe that the single-qubit operations and shuttling prefer a higher degree of parallelization than the two-qubit gates. Considering that xyD and zD have the same meaning, their combined averages (excluding -1 values) from the last row of Table 5 is 72.7% and the sD to 66.7%. This compares to an average of 52.6% for the two-qubit gates. Therefore, prioritizing the simultaneous operation of single-qubit gates (and shuttling) over two-qubit gates is more important in spin-qubit architectures.
@@
The swap_opt variable seems to show a preference for supporting SWAP gates for lower qubit counts. A possible explanation for this is that SWAP operations take a longer time than a parallelized sequence of shuttles [50]. As we explained before, the effects of having longer circuit execution times become more severe with larger circuits; hence, the extra SWAPs can outweigh the gained fidelity they initially offer. Combining this conclusion for large-scale circuits and the strong preferences for shuttling freedom (beSnake router) indicates that the shuttle operation is a key communication method for spin-qubit devices.
VIII Acknownledgement
@@
IX Conclusion
In this paper, we propose a comprehensive exploration of the current and future architectural design space for quantum dot spin-qubit quantum processors. Our research focuses on assessing the critical architectural characteristics that could be key in ensuring high performance on such devices. To this end, we present an upgraded version of the SpinQ compilation framework for spin qubit architectures in which architectural variables can configure internal compilation passes, making it the first Design Space Exploration (DSE) framework. After defining our design space, consisting of $29,312$ architectures, we propose a multi-optimization-based tool, ArtA (Artificial Architect), to automate the DSE process. The goal of ArtA is (a) to suggest which of the seventeen optimization method configurations can find the architecture with the highest ESP the fastest compared to the brute forcing approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per circuit category. We have shown up to 99.1% improvement in computation times compared to the brute-force approach, showing that it is possible to easily explore a vast number of designs. We have also provided insights into matching different optimization methods with specific quantum circuit categories. Bayesian optimization with an Expected Improvement acquisition function and a Matérn kernel was chosen as the best optimization method due to its effectiveness in managing higher qubit counts. After that, we proceed to automate the DSE process for 44 quantum circuits of up to 258 qubits within the 29,312-architecture design space. Our findings reveal the importance of minimizing the circuit duration by maximizing parallelization whenever possible instead of selecting architectural characteristics that minimize crosstalk errors. Crosstalk is currently a significant concern in experiments, but our long-term considerations deem decoherence more crucial. Lastly, the shuttle operation is preferred over costly SWAP operations for communication purposes, especially for large-scale circuits. Contrary to that, shuttle-based single-qubit gates such as Z rotations are not preferred over pulsed-based. Furthermore, our universal architecture suggestion places higher priority on parallelizing single-qubit gates rather than two-qubit gates. Overall, our work illustrates how the integration of DSE methodologies with optimization algorithms can be successfully utilized to offer valuable insights to designers of quantum processors.
As for future improvements, a broader range of hyper-parameter combinations and circuit variations could enhance ArtA’s effectiveness. Then, with advancements in spin qubit device fabrication, new architectural variables and values could potentially update the design space, and such integration will be relatively straightforward owing to the modularity of the internal functions of SpinQ and ArtA. Moreover, introducing non-architectural constraints such as development costs and time into ArtA’s design space could help make real-world conclusions regarding developing spin qubits. This would allow for operational fidelities to become variables as well and, consequently, add a second dimension to the Figure of Merit (ESP).
Also, some of the assumptions made in this work could be fine-tuned as the field progresses. It is crucial to acknowledge the impact of these assumptions on the observations presented. For instance, we assumed that the fidelity of pulse-based Z rotations is equivalent to that of a single shuttle operation. This assumption should not be construed as either favorable or unfavorable; rather, it is a foundational premise upon which ArtA bases its architectural “recommendations”. Should empirical evidence later demonstrate that shuttle operations exhibit significantly higher fidelity compared to pulse-based Z rotations or the crosstalk model depends on more factors, it may necessitate a reevaluation of preferences.
Further analysis of architectural patterns with similar quantum circuit characteristics can reveal more trends. For instance, quantum circuits with high two-qubit gate percentage ranges can perform best when the qubit connectivity of the architecture is above a certain degree. These kinds of statements or more can be made for a plethora of circuit characteristics [3, 51] and, in doing so, create a general matching between algorithm categories and architectural designs. Once this is done, it is possible to make the reversed suggestion of which architecture will perform best for an input circuit of interest once categorized without needing to repeat the DSE process.

### v3

@@ VII.2 Architecture analysis
Table 5: Architectures returned by BO [EI, Ma] optimization method for a range of quantum circuits (refer to Appendix X.1). Results are color-coded, with color schemes following the categories introduced in Section IV. Final-row averages omit -1 values.
@@
In Table 5, we present the values of the architectural variables, introduced in Section IV, that were found by ArtA having the highest ESP with the BO [EI, Ma] optimization method alongside their respected ESP values. These circuits are organized by class and number of qubits within the rows of the table. Subsequent columns present the ESP of the best architecture discovered, followed by color-coded architectural parameter categories as delineated in Section IV. The use of color accents in the table serves to enhance visual clarity, with each corresponding to the values of each variable.
@@
First, we observe a unanimous preference for the Z rotation implementation (z_rot_impl) in the zri column, with a pulse-based rotation being preferred. Note that previously, we have assumed in Section V.1 that a shuttle and a pulse-based Z rotation share the same single-qubit gate fidelity, but because the latter requires two shuttles, it can benefit the ESP more than the other type. Moreover, as a shuttle operation involves two quantum dots while a pulse-based operation only involves one, the shuttle-based Z rotation will have a higher chance of incurring crosstalk errors based on our crosstalk definition. Given these two reasons, BO has favored the pulse-based version. This indicates that shuttle-based single-qubit rotations are not sustainable compared to pulse-based ones for large-scale circuits as they progressively contribute more noise.
Moving on, since the X, Y, and Z gates are implemented the same way hardware-wise, xy_z automatically is assumed 1 based on the interdependencies detailed in Section IV.6. Consequently, xy_tqg and z_tqg are equivalent, taking both either 0 or 1 from the optimization method of ArtA (i.e., BO [EI, Ma] in this case), as we can observe in the table. Additionally, xy_z_tqg can also be conceptually equivalent to xy_tqg and z_tqg since pulse-based Z rotations are selected. Therefore, whenever xy_z_tqg is 1 or xy_tqg and z_tqg is 1, it means the architecture(s) which allow(s) parallelization of all single- and two-qubit gates obtains the highest ESP. In that sense, one would expect xy_z_tqg, xy_tqg and z_tqg to be synchronized. However, since such exploration freedom was given in the design space, the reasons for the mismatches could be explained by the internal characteristics of the circuits themselves. For example, in QFT100, the parallelization of all gate types occurs more often than in QFT60, as observed from their gate parallelization constraints xy_z_tqg = 0, xy_tqg = 1, and z_tqg = 1.
Conversely, when applying the same reasoning to QV8, QV10, and QFT20, it appears that hardware support for parallel execution of single- and two-qubit gates is not essential. This highlights that the success of real quantum circuits can stem from different architectural “angles,” even under reduced hardware capabilities [57]. This exemplifies the strength of our approach: it enables the identification of optimal architecture configurations without requiring the most complex or fabrication-intensive hardware designs.  [^3] Ultimately, each circuit profile has specific architectural traits that can be leveraged to achieve optimal performance without the need for excessively advanced hardware features (achieving a balance in architectural characteristics).
@@
The same can be observed in our universal architecture recommendation in the last row when rounding the numbers, which performs best on average for all the circuits examined. Additionally, the operational fidelity, which is directly linked to the gate overhead of compilation, can also severely impact the final ESP for circuits with high gate counts, which in turn explains the need to add the least additional gates possible during compilation through the use of beSnake [55] and local single-qubit implementation. Having said that, Bernstein Vazirani, QV8, and QV10 circuits preferred the global implementation, as these circuits consist of a significant number of single-qubit gates (with the same angles) being executed at the same time step.
@@
Overall, we can observe that the single-qubit operations and shuttling prefer a higher degree of parallelization than the two-qubit gates. Considering that xyD and zD have the same meaning due to zri = 1, their combined averages from the last row of Table 5 is 72.7%. This compares to an average of 52.6% for the two-qubit gates and 66.7% for sD. Therefore, prioritizing the simultaneous operation of single-qubit gates (and shuttling) over two-qubit gates is more important.
@@
The swap_opt variable seems to show a preference for supporting SWAP gates for lower qubit counts. A possible explanation for this is that SWAP operations take a longer time than a parallelized sequence of shuttles [55]. As we explained before, the effects of having longer circuit execution times become more severe with larger circuits; hence, the extra SWAPs can outweigh the gained fidelity they initially offer. Combining this conclusion for large-scale circuits with the observed strong preference for shuttling flexibility, enabled by the beSnake router, suggests that shuttle operations constitute a key communication mechanism in spin-qubit architectures for the future.
VIII Conclusion
In this paper, we propose a comprehensive exploration of the current and future architectural design space for quantum dot spin-qubit quantum processors. Our research focuses on assessing the critical architectural characteristics that could be key in ensuring high performance on such devices. To this end, we present an upgraded version of the SpinQ compilation framework [56] for spin qubit architectures in which architectural variables can configure internal compilation passes, resulting in the first Design Space Exploration (DSE) framework. After defining our design space, consisting of $29,312$ architectures, we propose a multi-optimization-based tool, ArtA (Artificial Architect), to automate the DSE process. The goal of ArtA is twofold: (a) to identify which of the seventeen optimization configurations can most efficiently discover the architecture that achieves the highest ESP, compared to brute-force exploration; and (b) to determine the architectural design characteristics that are critical for constructing high-performance spin-qubit devices across different circuit categories. We have shown up to 99.1% improvement in computation times compared to brute-forcing, showing that it is possible to easily explore a vast number of designs. We have also provided insights into matching different optimization methods with specific quantum circuit categories.
It should be noted, that the underlying goal of our study was not to exhaustively benchmark all possible optimization methods, but rather to introduce the concept of applying design space exploration to quantum computing systems – with a particular focus on spin-qubit architectures – by employing a diverse and representative selection of methods from different optimization families. Bayesian optimization with an Expected Improvement acquisition function and a Matérn kernel was chosen as the best method due to its effectiveness in managing higher qubit counts. After that, we proceed to automate the DSE process for 44 quantum circuits of up to 258 qubits within the 29,312-architecture design space. Our findings reveal the importance of minimizing the circuit duration by maximizing parallelization whenever possible instead of selecting architectural characteristics that minimize crosstalk errors. Crosstalk is currently a significant concern in experiments, but our long-term considerations deem decoherence more crucial. Lastly, the shuttle operation is preferred over costly SWAP operations for communication purposes, especially for large-scale circuits. Contrary to that, shuttle-based single-qubit gates such as Z rotations are not preferred over pulse-based. Furthermore, our universal architecture suggestion places higher priority on parallelizing single-qubit gates rather than two-qubit gates. Overall, our work demonstrates how the integration of DSE methodologies with optimization algorithms can uncover valuable architectural insights that would be difficult, if not impossible, to foresee while suggesting points of optimality with reduced hardware complexity.
However, it is important to acknowledge that while these findings highlight promising architectural directions, several of these design choices – such as high degrees of parallelization, connectivity, and flexible shuttling strategies – currently pose significant engineering challenges. Implementing fully local control, for instance, demands high fabrication uniformity and precise calibration, which is non-trivial at larger scales. Likewise, extensive shuttling flexibility requires precise individual gate addressing, which is experimentally demanding. Nonetheless, the purpose of this work is to chart a clear path forward, identifying architectural features that, while ambitious, can guide future efforts toward scalable and performant spin-qubit processors.
As for future improvements, a broader range of hyper-parameter combinations and circuit variations could enhance ArtA’s effectiveness. Then, with advancements in spin qubit device fabrication, new architectural possibilities could potentially expand the design space, and such integration will be relatively straightforward owing to the modularity of the internal functions of SpinQ and ArtA. Moreover, introducing non-architectural constraints such as development costs and time into ArtA’s design space could help make real-world predictions regarding developing spin qubits. For instance, this would allow for operational fidelities to become variables.
Also, some of the assumptions made in this work could be fine-tuned as the field progresses. It is crucial to acknowledge the impact of these assumptions on the observations presented. For instance, we assumed the fidelity of pulse-based Z rotations to be equivalent to that of a single shuttle operation. This assumption should not be construed as either favorable or unfavorable; rather, it is a foundational premise upon which ArtA bases its architectural “recommendations”. Should empirical evidence later demonstrate that shuttle operations exhibit significantly higher fidelity compared to pulse-based Z rotations or the crosstalk model depends on more factors, it may necessitate a reevaluation of preferences.
Further analysis of architectural patterns associated with specific quantum circuit characteristics may reveal additional trends and enable systematic categorization. For example, circuits with a high proportion of two-qubit gates (e.g., 90% or more) may exhibit optimal performance only when the qubit connectivity exceeds a certain threshold. Insights like these, based on various circuit properties [3, 56], can inform the construction of a generalized lookup table that maps algorithmic profiles to architectural configurations. Such a resource would enable rapid prediction of high-performing architectures for new circuits, eliminating the need to re-run the full DSE process.
IX Acknownledgement
@@
Declarations
There are no competing interests between the authors.